OpenStudy (anonymous):

Can someone please help me solve y = f(x)= x-3/2x+4 im trying to find it's inverse

OpenStudy (anonymous):

ok is this your question?? \[\Large y=\frac{x-3}{2x+4}\]

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

ok first of all replace all x with y's and all y's with x \[\Large x=\frac{y-3}{2y+4}\] is it ok with you ?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

now solve the above for y. can you do this ?

OpenStudy (anonymous):

That's were I get confused, I'm not sure on how to eliminate the format as a fraction

OpenStudy (anonymous):

sorry openstudy crashed me :( let me type again.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ok can i write this as \[\Large x(2y+4)=y-3\] ?

OpenStudy (anonymous):

sure, I did that. Then would I make it 2xy+4x = y-3?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

then divide the y out?

OpenStudy (anonymous):

now can i take y to the left side and 4x to the right \[\Large 2xy-y=-3-4x\] it it ok with you ?

OpenStudy (anonymous):

is*

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

ok now taking y common from the left side \[\Large \implies (2x-1)y=-3-4x\] now ok ??

OpenStudy (anonymous):

I'm sorry, I'm a bit confused

OpenStudy (anonymous):

ok take common y from the following and tell me what you got \[\Large 2xy-y=-3-4x\]

OpenStudy (anonymous):

it would be (2x-1)y = -3-4x

OpenStudy (anonymous):

yes exactly!

OpenStudy (anonymous):

I know that I should try to get all my x's on one side, so I would I divide both sides by 2x-1

OpenStudy (anonymous):

well done that's true !!!!!

OpenStudy (anonymous):

Do I stop there, or continue?

OpenStudy (anonymous):

Just one more step !! after division by 2x-1 you get \[\Large y=\frac{-3-4x}{2x-1}\] agree?

OpenStudy (anonymous):

Yes, If I continue to do so... I would get y = 3-2x?

OpenStudy (anonymous):

I hope I'm not making my own math rules

OpenStudy (anonymous):

I mean y = 3+2x

OpenStudy (anonymous):

so now you have \[\Large y=\frac{-3-4x}{2x-1}\] just replace y with \[ f^{-1}(x)\] and it is your answer \[\Large \implies f^{-1}(x)=\frac{-3-4x}{2x-1}\]

OpenStudy (anonymous):

ok, thanks. How would I graph that?

OpenStudy (anonymous):

graphing is a whole different story !!!

OpenStudy (anonymous):

just to give you hint take different values of x .put in the above expression. you will get corresponding values of f^(-1)(x) Plot x values on x axis and f^-1(x) values on y axis.